# Emergent Symbolic Mechanisms Support

### v1

@@ Abstract
Many recent studies have found evidence for emergent reasoning capabilities in large language models, but debate persists concerning the robustness of these capabilities, and the extent to which they depend on structured reasoning mechanisms. To shed light on these issues, we perform a comprehensive study of the internal mechanisms that support abstract rule induction in an open-source language model (Llama3-70B). We identify an emergent symbolic architecture that implements abstract reasoning via a series of three computations. In early layers, symbol abstraction heads convert input tokens to abstract variables based on the relations between those tokens. In intermediate layers, symbolic induction heads perform sequence induction over these abstract variables. Finally, in later layers, retrieval heads predict the next token by retrieving the value associated with the predicted abstract variable. These results point toward a resolution of the longstanding debate between symbolic and neural network approaches, suggesting that emergent reasoning in neural networks depends on the emergence of symbolic mechanisms.
@@ 1 Introduction
Large language models have become the dominant paradigm in artificial intelligence, but there is significant ongoing debate concerning the limits and reliability of their capabilities. One major focus of this debate has been the question of whether they can reason systematically in a abstract or human-like manner. Many studies have documented impressive performance on various reasoning tasks (Wei et al., 2022; Mirchandani et al., 2023), even rivaling human performance in some cases (Webb et al., 2023; Musker et al., 2024; Webb et al., 2024a), but other studies have questioned these conclusions (Wu et al., 2023; McCoy et al., 2023; Lewis & Mitchell, 2024). In particular, language models appear to perform more poorly in some reasoning domains, such as mathematical reasoning (Dziri et al., 2024) or planning (Momennejad et al., 2024); and, even in domains in which they have shown strong performance such as analogical reasoning (Webb et al., 2023), some studies have questioned the robustness of these capabilities (Lewis & Mitchell, 2024).
@@
These conflicting findings raise the question: are language models genuinely capable of abstractly structured human-like reasoning, or are they merely mimicking this capacity by statistically approximating their training data? One way to answer this question is to look at the internal mechanisms that support this capacity. It has long been hypothesized that innate symbol-processing mechanisms are required to support human-like abstraction (Marcus, 2001; Dehaene et al., 2022; Wong et al., 2023). It has also been demonstrated that neural networks are capable, at least in principle, of implementing some of the key properties of symbolic systems (Smolensky, 1990; Hummel & Holyoak, 2003; Kriete et al., 2013), and that the incorporation of these properties as architectural inductive biases can support data-efficient acquisition of abstract symbolic reasoning (Webb et al., 2020; Altabaa et al., 2023; Webb et al., 2024b). It remains unclear, however, in the case of transformer language models that do not obviously possess such strong inductive biases, what mechanisms support their emergent capability for abstraction.
@@
Figure 1: Emergent Symbolic Architecture. Schematic depiction of the proposed three-stage architecture for abstract reasoning in language models. Inputs are identity rule problems, each instance of which is constructed from arbitrarily chosen tokens (example shown is an ABA rule). Symbol abstraction heads identify relations between input tokens and, based on these relations, represent the tokens using a consistent set of abstract variables aligned with their role in the relations. Symbolic induction heads perform sequence induction over abstract variables (i.e., they predict the next variable based on the sequence observed in the previous in-context examples). Retrieval heads predict the next token by retrieving the value associated with the predicted abstract variable.
Here, we report evidence for a set of emergent symbolic mechanisms that support abstract reasoning in large language models. We focus on a simple but paradigmatic abstract reasoning task–induction of algebraic rules–and identify an emergent three-stage architecture that supports performance of this task in the open-source language model Llama3-70B (Dubey et al., 2024). Notably, this architecture can be straightforwardly interpreted as performing a form of symbol processing. Specifically, we find evidence to support a three-stage procedure in which: 1) input tokens are converted to abstract variables (i.e., symbols) based on the relations between those tokens, 2) sequence induction is then performed over these variables, and 3) next-token prediction is performed by retrieving the value associated with the predicted variable. We refer to the attention heads that perform these computations as symbol abstraction heads, symbolic induction heads, and retrieval heads respectively. Through a series of ablation and causal mediation experiments, we show that this architecture is both necessary and sufficient to perform abstract rule induction. We also carry out representational similarity analyses and analyze attention patterns to better understand these mechanisms, finding evidence that they operate over symbol-like representations. Taken together, these results suggest that emergent reasoning in large language models depends on structured, abstract reasoning mechanisms, rather than less abstract forms of statistical approximation. More broadly, these results suggest a resolution to the longstanding debate between symbolic and neural network approaches, illustrating how neural networks can learn to perform abstract reasoning via the development of emergent symbol processing mechanisms.
@@ 2 Approach
Figure 1 depicts the proposed architecture and rule induction task. In this work, we focused on an algebraic rule induction task involving sequences governed by one of two identity rules, ABA or ABB. For each problem, two in-context examples were presented, followed by an incomplete third example. The model was expected to generate the token that completes this third example. We instantiated rules using tokens randomly sampled from Llama3’s vocabulary (ensuring that in-context examples within the same problem instance did not share tokens). We found that Llama3-70B displayed a 2-shot accuracy of 95% on this task.
Although this task is relatively simple, especially when compared with some of the tasks that have been featured in recent debates over LLM reasoning (e.g., matrix reasoning tasks (Webb et al., 2023)), it nevertheless offers a paradigmatic case of relational abstraction. In particular, the use of completely arbitrary tokens ensures that the task cannot be solved by relying on statistical patterns specific to the tokens or associations among them, and for this reason it has previously been used to argue for the presence of symbol-processing mechanisms in human cognition (Marcus et al., 1999), and to evaluate systematic generalization of abstract rules in neural networks (Webb et al., 2020). Accordingly, the ability to reliably solve this task is already strongly suggestive of the presence of some form of symbol-processing. In the following sections, we describe a specific mechanistic hypothesis for how symbol-processing might be carried out in this model.
@@ 2.1 Symbol Abstraction Heads
Our hypothesis consists of three stages. In the first stage, input tokens are converted to symbolic representations. The inspiration for this hypothesis comes from the abstractor architecture (Altabaa et al., 2023), a variant of the transformer that implements a strong relational inductive bias (Webb et al., 2024b). In that architecture, a modified form of attention (termed relational cross-attention) is employed in which the values consist of a standalone set of learned embeddings, rather than being conditioned on the input tokens as in standard self attention. As a result, the output of this attention operation is completely abstracted away from the identity of the input tokens, and instead only reflects the pattern of relations among those tokens (as encoded by the pattern of inner products between query and key embeddings). These outputs can therefore be viewed as a form of learned, distributed symbolic representations.
Here, we hypothesize that an emergent form of this relational attention operation is implemented by attention heads in early layers of the model. We refer to these heads as symbol abstraction heads. Concretely, the keys and query embeddings in these heads represent the input tokens, and the inner product between keys and queries represents the relations between these tokens. It is natural to interpret this operation as representing similarity relations (and this is the relevant type of relations in our task), but it is also possible for this operation to represent a broader class of relations (Altabaa & Lafferty, 2024). Importantly, we hypothesize that the value embeddings in these heads do not carry information about the specific identity of the input tokens, but instead represent only their position. More precisely, we hypothesize that the value embeddings represent the relative position of a token within an in-context example, as this is precisely the information that’s needed to compute the abstract variable associated with that token (e.g., the fact that the first token and the third token are the same in an ABA rule is precisely what determines that they share the same variable). Given that these conditions are met, the self-attention operation is equivalent to relational cross-attention (Altabaa et al., 2023), and the output of such an attention head will be analogous to an abstract variable.
@@ 2.2 Symbolic Induction Heads
In the second stage, we hypothesize that sequence induction is performed over the abstract variables computed in the first stage. This hypothesis is inspired by previous work on induction heads, an emergent circuit that supports in-context learning in transformers (Elhage et al., 2021; Olsson et al., 2022). As originally formulated, this circuit performs a simple sequence induction mechanism: given a sequence that ends with a particular token, an induction head will look for previous instances of that token, and retrieve the token that succeeded it. Although this mechanism performs induction based only on in-context bigram statistics, subsequent work has identified heads that also compute higher-order n-gram statistics (Akyürek et al., 2024). Here, we use the term ‘induction’ to refer to the more general process of predicting the next token based on in-context transition probabilities (i.e., beyond bigram statistics).
@@ 2.3 Retrieval Heads
Finally, in the third stage, we hypothesize that a separate mechanism is used to convert the abstract variables (symbols) to their associated tokens (values), by performing a simple form of retrieval. We refer to the attention heads that perform this retrieval operation as retrieval heads. The key and query embeddings in these heads represent abstract variables, and the value embeddings represent the corresponding input tokens. Retrieval heads perform the inverse of the relational attention operation performed by symbol abstraction heads. Given an input embedding representing an abstract variable (the prediction computed by symbolic induction heads in previous layers), this variable is matched with previous instances, and the associated token is retrieved. This can be viewed as a form of the ‘dereferencing’ operation that is central to symbolic computing, wherein a variable (i.e., a pointer to a particular location in memory) is used to retrieve the value associated with it (i.e., the data stored at that location).
@@ 3.1 Causal Mediation Analyses
We performed causal mediation analysis (Pearl, 2022; Meng et al., 2022; Wang et al., 2022; Todd et al., 2023) to isolate the hypothesized attention heads. In this analysis, embeddings from one context are patched into another context. This approach can be used to estimate the causal effect of an embedding at a particular layer, position, or attention head.
Our analysis had two conditions. In one condition, intended to isolate representations of abstract variables (i.e., symbols), we created two contexts in which the same token is associated with two different variables. Given a set of tokens $A_{1},B_{1},...A_{N},B_{N}$ (where $N-1$ represents the number of in-context examples), we created one context $c_{1}^{abstract}$ that instantiated an ABA rule, and another context $c_{2}^{abstract}$ that instantiated an ABB rule:
@@
In this condition, both contexts involve the same abstract rule, but the tokens used in the final example ($A_{n}$ vs. $B_{n}$) are swapped. We also perform a version of both analyses using the ABB rule and average the results for the two rules.
Together, these analyses allow for a double dissociation between representations of abstract variables (abstract causal mediation) vs. the tokens associated with those variables (token causal mediation). For each analysis, given two contexts $c_{1}$ and $c_{2}$, and a pretrained lanuage model $f(c)$ that outputs logits for all possible next tokens, we computed the causal mediation score developed by Wang et al. (2022):
$s=(f(c_{1}^{*})[y_{c_{1}^{*}}]-f(c_{1}^{*})[y_{c_{1}}])-(f({c_{1}})[y_{c_{1}^{*}}]-f({c_{1}})[y_{c_{1}}]),$ (5)
where $f(c_{1})[y]$ is the logit for answer $y$ in context $c_{1}$; $f(c_{1}^{*})[y]$ is the logit for answer $y$ in the patched context $c_{1}^{*}$, for which activations are patched from $c_{2}$ to $c_{1}$ at specific layers, positions, and/or attention heads; $y_{c_{1}}$ is the correct answer for context $c_{1}$; and $y_{c_{1}^{*}}$ is the expected answer for the patched context $c_{1}^{*}$. We averaged over different sets of tokens. The procedure for causal mediation analyses is shown in Algorithm 1. Intuitively, this score reflects the extent to which patching activations from context $c_{2}$ to $c_{1}$ has the expected effects on the model’s outputs.
@@ 3.2 Attention Analyses
Figure LABEL:symb_ind_heads_attention shows the results for symbolic induction heads. Our hypothesis predicts that attention should be directed from the final position in the entire sequence (the separation token at the end of the incomplete example) to positions of the final items in each in-context example, as these are the positions where the previous instances of the to-be-predicted abstract variable are located. The results confirmed this hypothesis. Again, the pattern was stronger for the second in-context example, consistent with a general effect of in-context learning, and consistent with the more focused pattern of attention observed for the symbol abstraction heads in the second in-context example.
Figure LABEL:retrieval_heads_attention) shows the results for retrieval heads. Our hypothesis predicts that attention should be directed from the final position in the sequence to the positions corresponding to the tokens that will appear next (i.e., within the incomplete example). For ABA rules (top), we predict that attention should be directed to the first item in the example (corresponding to the variable A), and for ABB rules (bottom) we predict that attention should be directed to the second item in each example (corresponding to the variable B). This prediction too was confirmed by our analyses.

### v2

@@ Abstract
Many recent studies have found evidence for emergent reasoning capabilities in large language models (LLMs), but debate persists concerning the robustness of these capabilities, and the extent to which they depend on structured reasoning mechanisms. To shed light on these issues, we study the internal mechanisms that support abstract reasoning in LLMs. We identify an emergent symbolic architecture that implements abstract reasoning via a series of three computations. In early layers, symbol abstraction heads convert input tokens to abstract variables based on the relations between those tokens. In intermediate layers, symbolic induction heads perform sequence induction over these abstract variables. Finally, in later layers, retrieval heads predict the next token by retrieving the value associated with the predicted abstract variable. These results point toward a resolution of the longstanding debate between symbolic and neural network approaches, suggesting that emergent reasoning in neural networks depends on the emergence of symbolic mechanisms.
@@ 1 Introduction
Large language models (LLMs) have become the dominant paradigm in artificial intelligence, but there is ongoing debate concerning the limits and reliability of their capabilities. One major focus of this debate has been the question of whether they can reason systematically in an abstract or human-like manner. Many studies have documented impressive performance on various reasoning tasks (Wei et al., 2022; Mirchandani et al., 2023), even rivaling human performance in some cases (Webb et al., 2023; Musker et al., 2024; Webb et al., 2025), but other studies have questioned these conclusions (Wu et al., 2024; McCoy et al., 2023; Lewis & Mitchell, 2024). In particular, LLMs appear to perform more poorly in some reasoning domains, such as mathematical reasoning (Dziri et al., 2024) or planning (Momennejad et al., 2024); and, even in domains in which they have shown strong performance such as analogical reasoning (Webb et al., 2023), some studies have questioned the robustness of these capabilities (Lewis & Mitchell, 2024).
@@
These conflicting findings raise the question: are LLMs genuinely capable of structured human-like reasoning, or are they merely mimicking this capacity by statistically approximating their training data? One way to answer this question is to look at the internal mechanisms that support this capacity. It has long been hypothesized that innate symbol-processing mechanisms are required to support human-like abstraction (Marcus, 2001; Dehaene et al., 2022; Wong et al., 2023). It has also been demonstrated that neural networks are capable, at least in principle, of implementing some of the key properties of symbolic systems (Smolensky, 1990; Hummel & Holyoak, 2003; Kriete et al., 2013), and that the incorporation of these properties as architectural inductive biases can support data-efficient acquisition of abstract symbolic reasoning (Webb et al., 2021; Altabaa et al., 2024; Webb et al., 2024). It remains unclear, however, in the case of transformer-based LLMs that do not obviously possess such strong inductive biases, what mechanisms support their emergent capability for abstraction.
@@
Figure 1: Emergent Symbolic Architecture. Schematic depiction of the proposed three-stage architecture for abstract reasoning in language models. Example depicts architecture as applied to an algebraic rule induction task, involving either ABA or ABB rules instantiated with random tokens. Symbol abstraction heads identify relations between input tokens and, based on these relations, represent the tokens using a consistent set of abstract variables aligned with their role in the relations. Symbolic induction heads perform sequence induction over abstract variables (i.e., they predict the next variable based on the sequence observed in the previous in-context examples). Retrieval heads predict the next token by retrieving the value associated with the predicted abstract variable.
Here, we report evidence for a set of emergent symbolic mechanisms that support abstract reasoning in LLMs. Specifically, we identify an emergent three-stage architecture consisting of the following mechanisms: 1) in early layers, symbol abstraction heads convert input tokens to abstract variables (i.e., symbols) based on the relations between those tokens, 2) in intermediate layers, symbolic induction heads, perform sequence induction over these variables, and 3) in later layers, retrieval heads perform next-token prediction by retrieving the value associated with the predicted variable. These mechanisms capture two key properties of symbol processing. First, they operate over representations of abstract variables that are invariant to the values that they are associated with. That is, the representation of an abstract variable is the same regardless of which values instantiate that variable. Second, these mechanisms utilize indirection, meaning that variables refer to content that is stored at a different location than the variables themselves (i.e., they are pointers).
To identify and validate the presence of these mechanisms, we perform experiments across three abstract reasoning tasks – algebraic rule induction, letter string analogies, and verbal analogies – and 13 open-source LLMs from four model families – GPT-2 (Radford et al., 2019), Gemma-2 (Gemma Team, 2024), Qwen2.5 (Qwen Team, 2025), and Llama-3.1 (Dubey et al., 2024) – drawing on convergent evidence from a series of causal, representational, and attention analyses. We find robust evidence for these mechanisms across all three tasks, and three out of four model families (Gemma-2, Qwen2.5, and Llama-3.1; with more equivocal results for GPT-2). Taken together, these results suggest that emergent reasoning in LLMs depends on structured, abstract reasoning mechanisms, rather than simple statistical approximation. More broadly, these results suggest a resolution to the longstanding debate between symbolic and neural network approaches, illustrating how neural networks can learn to perform abstract reasoning via the development of emergent symbol processing mechanisms.
@@ 2 Approach
Figure 1 depicts the proposed architecture as applied to an algebraic rule induction task involving sequences governed by one of two identity rules, ABA or ABB. For each problem, two in-context examples were presented, followed by an incomplete third example. The model was expected to generate the token that completes this third example. We instantiated rules using tokens randomly sampled from each model’s vocabulary (ensuring that in-context examples within the same problem instance did not share tokens). We begin by presenting results from this rule induction task and the Llama-3.1 70B model, and then present results for additional tasks and models.
We found that Llama-3.1 70B displayed a 2-shot accuracy of 95% on the rule induction task. Although this task is relatively simple, especially when compared with some of the tasks that have been featured in recent debates over LLM reasoning (e.g., matrix reasoning (Webb et al., 2023)), it nevertheless offers a paradigmatic case of relational abstraction. In particular, the use of completely arbitrary tokens ensures that the task cannot be solved by relying on statistical patterns specific to the tokens or associations among them, and for this reason it has previously been used to argue for the presence of symbol-processing mechanisms in human cognition (Marcus et al., 1999), and to evaluate systematic generalization of abstract rules in neural networks (Webb et al., 2021). Accordingly, the ability to reliably solve this task is already strongly suggestive of the presence of some form of symbol-processing. In the following sections, we describe a specific mechanistic hypothesis for how symbol-processing might be carried out in this model.
@@ 2.1 Symbol Abstraction Heads
Our hypothesis consists of three stages. In the first stage, input tokens are converted to symbolic representations. The inspiration for this hypothesis comes from the abstractor architecture (Altabaa et al., 2024), a variant of the transformer that implements a strong relational inductive bias (Webb et al., 2024). In that architecture, a modified form of attention (termed relational cross-attention) is employed in which the values consist of a standalone set of learned embeddings, rather than being conditioned on the input tokens as in standard self-attention. As a result, the output of this attention operation is completely abstracted away from the identity of the input tokens, and instead only reflects the pattern of relations among those tokens (as encoded by the pattern of inner products between query and key embeddings). These outputs can therefore be viewed as a form of learned, distributed symbolic representations.
Here, we hypothesize that an emergent form of this relational attention operation is implemented by attention heads in early layers of the model. We refer to these heads as symbol abstraction heads. Concretely, the keys and query embeddings in these heads represent the input tokens, and the inner product between keys and queries represents the relations between these tokens. It is natural to interpret this operation as representing similarity relations (and this is the relevant type of relations in the rule induction task), but it is also possible for this operation to represent a broader class of relations (Altabaa & Lafferty, 2024) (as necessitated by some of the other tasks that we investigate). Importantly, we hypothesize that the value embeddings in these heads do not carry information about the specific identity of the input tokens (i.e., they are invariant to the content of those tokens), but instead represent only their position. More precisely, we hypothesize that the value embeddings represent the relative position of a token within an in-context example, as this is precisely the information that’s needed to compute the abstract variable associated with that token (e.g., the fact that the first token and the third token are the same in an ABA rule is precisely what determines that they share the same variable). Given that these conditions are met, the self-attention operation is equivalent to relational cross-attention (Altabaa et al., 2024), and the output of such an attention head will represent an abstract variable.
@@ 2.2 Symbolic Induction Heads
In the second stage, we hypothesize that sequence induction is performed over the abstract variables computed in the first stage. This hypothesis is inspired by previous work on induction heads, an emergent circuit that supports in-context learning in transformers (Elhage et al., 2021; Olsson et al., 2022). As originally formulated, this circuit performs a simple sequence induction mechanism: given a sequence that ends with a particular token, an induction head will look for previous instances of that token, and retrieve the token that succeeded it. Although this mechanism performs induction based only on in-context bigram statistics, subsequent work has identified heads that also compute more complex n-gram statistics (Akyürek et al., 2024). Here, we use the term ‘induction’ to refer to the more general process of predicting the next token based on in-context transition probabilities (i.e., beyond bigram statistics).
@@ 2.3 Retrieval Heads
Finally, in the third stage, we hypothesize that a separate mechanism is used to convert the abstract variables (symbols) to their associated tokens (values), by performing a simple form of retrieval. We refer to the attention heads that perform this retrieval operation as retrieval heads. The key and query embeddings in these heads represent abstract variables, and the value embeddings represent the corresponding input tokens. Retrieval heads perform the inverse of the relational attention operation performed by symbol abstraction heads. Given an input embedding representing an abstract variable (the prediction computed by symbolic induction heads in previous layers), this variable is matched with previous instances, and the associated token is retrieved. This can be viewed as a form of indirection, wherein a variable (i.e., a pointer to a particular location in memory) is used to retrieve the value associated with it (i.e., the data stored at that location).
2.4 Binding
It is worth noting that our hypothesized mechanisms depend on the assumption that LLMs can perform variable binding, i.e., that they can represent in-context associations between variables and their values. For instance, as depicted in Figure 1, the retrieval head mechanism depends on the assumption that the model has bound the predicted abstract variable (A) to its associated token (‘ks’), enabling the variable to be used as a pointer to retrieve the token. This capacity for variable-binding is another key component of symbol processing.
Previous work has found that transformer-based LLMs implement variable-binding via a binding ID mechanism (Feng & Steinhardt, 2024). Binding IDs are vectors that represent indices used for keeping track of in-context associations (e.g., for binding entities to their attributes). Prevous work has suggested that binding IDs are additively combined with the embeddings for the entities and the attributes that they represent (Feng & Steinhardt, 2024). We assume that a similar mechanism supports variable-binding in the context of our hypothesized architecture, i.e., that representations of abstract variables are bound to their associated tokens by adding them to the residual stream at the position of those tokens. It should also be noted that this binding mechanism depends on the presence of discrete slots (e.g., separate embeddings for each input token), a feature that is shared by both the transformer architecture (Vaswani et al., 2017) and other external memory architectures (Graves et al., 2016).
@@ 3.1 Causal Mediation Analyses
We performed causal mediation analysis (Pearl, 2022; Meng et al., 2022; Wang et al., 2023; Todd et al., 2024) to isolate the hypothesized attention heads. In this analysis, embeddings from one context are patched into another context. This approach can be used to estimate the causal effect of an embedding at a particular layer, position, or attention head.
Our analysis had two conditions. In one condition, intended to isolate representations of abstract variables (i.e., symbols), we created two contexts in which the same token is associated with two different variables. Given a set of tokens $A_{1},B_{1},...A_{N},B_{N}$ (where $N-1$ represents the number of in-context examples), we created one context $c_{1}^{abstract}$ that instantiated an ABA rule, and another context $c_{2}^{abstract}$ that instantiated an ABB rule (equivalent to a BAA rule):
@@
In this condition, both contexts involve the same abstract rule, but the tokens used in the final example are swapped. We also perform a version of both analyses using the ABB rule and average the results for the two rules.
Together, these analyses allow for a double dissociation between representations of abstract variables (abstract causal mediation) vs. the tokens associated with those variables (token causal mediation). For each analysis, given two contexts $c_{1}$ and $c_{2}$, and a pretrained lanuage model $f(\cdot)$ that outputs logits for all possible next tokens, we computed the causal mediation score developed by Wang et al. (2023):
$s$ $=\left(f(c_{1}^{*})[y_{c_{1}^{*}}]-f(c_{1}^{*})[y_{c_{1}}]\right)$
$\quad-\left(f(c_{1})[y_{c_{1}^{*}}]-f(c_{1})[y_{c_{1}}]\right)$ (5)
where $f(c_{1})[y]$ is the logit for answer $y$ in context $c_{1}$; $f(c_{1}^{*})[y]$ is the logit for answer $y$ in the patched context $c_{1}^{*}$, for which activations are patched from $c_{2}$ to $c_{1}$ at specific layers, positions, and/or attention heads; $y_{c_{1}}$ is the correct answer for context $c_{1}$; and $y_{c_{1}^{*}}$ is the expected answer for the patched context $c_{1}^{*}$. Intuitively, this score reflects the extent to which patching activations from context $c_{2}$ to $c_{1}$ has the expected effects on the model’s outputs. For $c^{abstract}$, which was designed to test whether the activations represent abstract variables, we expected that activation patching would convert the symbol for the token that completes the incomplete query, whereas for $c^{token}$, we expected that activation patching would convert the literal token. See Section B.2 and Figures 6-8 for more details (including illustrated examples) on this analysis.
@@ 3.2 Attention Analyses
Figure LABEL:symb_ind_heads_attention shows the results for symbolic induction heads. Our hypothesis predicts that attention should be directed from the final position in the entire sequence (the separation token at the end of the incomplete example) to positions of the final items in each in-context example, as these are the positions where the previous instances of the to-be-predicted abstract variable are located. The results confirmed this hypothesis. Again, the pattern was stronger for the second in-context example, consistent with a general effect of in-context learning.
Figure LABEL:retrieval_heads_attention) shows the results for retrieval heads. Our hypothesis predicts that attention should be directed from the final position in the sequence to the positions corresponding to the tokens that will appear next (i.e., within the incomplete example). For ABA rules (top), we predict that attention should be directed to the first item in the example (corresponding to the variable A), and for ABB rules (bottom) we predict that attention should be directed to the second item in each example (corresponding to the variable B). This prediction too was confirmed by our analyses. Taken together, these results suggest that the model employs a form of indirection, first computing the variable associated with the predicted token, and then using that variable as a pointer to retrieve the token itself.